Chapter 27, Problem 03 Your answer is partially correct. Try again. Two in-phase sources of waves are separated by a distanceWhat is b and c? Please show work.
Answer #1

___________________________

(a)

destructive

___________________________

(b)

distance between two sources, d = 3.94 m

let x is the distance from one source where destructive interference takes place.

so, path difference, r2 - r1 = lamda/2

(d-x) - x = lamda/2

3.94 - x - x = 5.30/2

3.94 - 2*x = 2.65

2*x = 3.94-2.65

x = (3.94-2.65)/2

x = 0.645 m  

___________________________

(c)

d - x = 3.94 - 0.645

d = 3.295 m

___________________________

Please rate

if any mistake in this answer please comment i will clarify your doubt . thank you